how can we know that equation has no solution?im a bit confusing about that.
check this out ; x=rootx-6 ,ive been work it out and find that x=3 x=-2 then i look at the answers from the back and it just written as no solution .really confused! -_-

Answer by stanbon(75887)

how can we know that equation has no solution?
I'm a bit confused about that.
check this out ; x = root(x-6).
I've been working it out and find that x=3 x=-2 then i look at the answers from the back.It says no solution. Really confused!
----------------
Your Problem:
x = sqrt(x-6)
Square both sides to get:
x ^2 = x-6
x^2 - x + 6 = 0
(x-3)(x+2) = 0
x = 3 or x = -2
-----
Checking the answers in the original equation.
Can x = 3 ?
No because sqrt(3-6) = sqrt(-3) is not a Real Number.
--------------
Can x = -2 ?
No because sqrt(-2-6) = sqrt(-8) is not a Real Number.
---------------
Therefore there is no solution for your problem.
===============================================
Cheers,
Stan H.